OpenStudy (anonymous):

Can someone please help? :c Using complete sentences, explain how to find the maximum value for each function and determine which function has the largest maximum y-value. f(x)=-4(x-6)^2 +3. g(x)=2cos(2x-pi)+4

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@ganeshie8 @phi can either of you help?

OpenStudy (anonymous):

*dies*

OpenStudy (phi):

is this calculus? if so, take the derivative of the function, and set to 0 to find where it has a max (or min). However, you do not need calculus. The first function, f(x)=-4(x-6)^2 + 3 is a parabola ( in the shape of a "frown" ) with a max value at its vertex. because they gave you the equation in "vertex form" y = a(x-h)^2 + k , you can read off the vertex to be (h,k) In this case, the vertex is at (6,3) for the second function, g(x)=2cos(2x-pi)+4 you should know that the max value of the cosine is 1, so the max will be g(x)= 2*1 + 4 = 6